Grade 1010 grade maths

Use euclide division algorithm to find hcf of 129 and 545 and hence express it in the form of 129x +545y

Migaya Benhur
9 Years agoGrade 10

5 Answers

Migaya Benhur
9 Years ago
hcf of 129 and 545 to find 129x+545y
545=129*4+29
129=29*4+13
29=13*2+3
13=3*4+1
3=1*3+0
 
29=545/4-129
 
y=1/4
x=1
this is the answer
sachin
9 Years ago
 
hcf of 129 and 545 to find 129x+545y
545=129*4+29
129=29*4+13
29=13*2+3
13=3*4+1
3=1*3+0
 
29=545/4-129
 
y=1/4
x=1
sachin
9 Years ago
 
hcf of 129 and 545 to find 129x+545y
545=129*4+29
129=29*4+13
29=13*2+3
13=3*4+1
3=1*3+0
 
29=545/4-129
 
y=1/4
x=1.
sachin
9 Years ago
 
hcf of 129 and 545 to find 129x+545y
545=129*4+29
129=29*4+13
29=13*2+3
13=3*4+1
3=1*3+0
 
29=545/4-129
 
y=1/4
x=only1
sachin
9 Years ago
hcf of 129 and 545 to find 129x+545y
545=129*4+29
129=29*4+13
29=13*2+3
13=3*4+1
3=1*3+0
 
29=545/4-129
 
y=1/4
x=1